A young boy swings a yo-yo horizontally above his head so that the yo-yo has a cen- tripetal acceleration of 250 m/s2.
If the yo-yo’s string is 0.14 m long, what is the yo-yo’s tangential speed?
Answer in units of m/s
1
Expert's answer
2011-10-27T11:18:16-0400
Tangential speed Vt=omega * R

where R=0.14 m - the yo-yo’s string
omega - angular velocity of the yo-yo

centripetal acceleration Wc=omega^2 *R
=> omega=sqrt(Wc/R)=sqrt(250/0.14)=42.25
Hence Vt=omega*R=42.25*0.14=5.91 m/s